How much 96% ethanol do you need to mix with water to prepare 100ml of 60% ethanol? Please show working
Answer #1

You can use the equation shown below:

C1 x V1 = C2 x V2

This says that the concentration you begin with, multiplied by the volume you use of it, will be equal to the final concentration, multiplied by whatever the final volume is.


C1 = the strength of the alcohol that we should have
V1 = the volume of the alcohol that we should have
C2 = the strength of the alcohol that we will use instead
V2 = the volume of the alcohol that we will use instead

Calculation:

C1= 96% or 0.96, V1=?, C2= 60% or 0.60, V2= 100

C1xV1=C2xV2

V1=\frac{C2 \times V2}{C1}


V1=\frac{0.60 \times 100}{0.96} =62.5

Thus, 62.5ml of 96% of ethanol should be added to water to make 100ml of 60% ethanol.